# On the concentration distribution in turbulent thermals

### v1

@@ 3.2 Numerical simulations
We have conducted 3D simulations of thermals with the open-source pseudospectral code Dedalus (Burns et al., 2016, 2020), using the setup described by Lecoanet & Jeevanjee (2019). The equations of conservation of momentum and concentration are solved under the Boussinesq approximation. The thermal is initialized at $t=0$ by introducing a spherical concentration perturbation of diameter $d$ into an otherwise homogeneous fluid. The spatial domain extends horizontally over $10d$ in both the $x$ and $y$ directions, and vertically over $20D$ in the $z$ direction. It is respectively discretized in $256\times 256\times 512$ modes. We benchmark our simulations against a simulation with a higher resolution of $512\times 512\times 1024$, kindly provided by Daniel Lecoanet (Lecoanet & Jeevanjee, 2019). At a given Reynolds number ($Re\approx 4631$), we find that the kinetic energy spectrum and PDF of concentration (figure 23) are very similar between the two resolutions.
@@
Our numerical simulations use Reynolds numbers that match those of our experiments. In contrast, the Schmidt number $Sc$ is set to 1 in all numerical simulations (for computational resource reasons), resulting in a concentration field that diffuses much faster than in the experiments, where $Sc=2500$. Consequently, the Péclet number is also a factor $2500$ larger in the experiments than in the simulations. It will allow us to investigate the effect of the Schmidt number on concentration distributions.
@@
In each simulation, we extract two orthogonal planes that intersect in the middle of the grid. In Section 5, we will show the evolution of concentration moments and PDF for the average of these two orthogonal planes and compare them to the experiments performed at the same Reynolds numbers.
4 Qualitative description of a turbulent thermal
4.1 Macroscopic evolution
@@
Figure 2 shows the evolution of the concentration field for a given $Re=3452$ for one experiment and a simulation. The concentration scale changes between each time step for better visualization. During its fall, the radius of the turbulent thermal increases as the surrounding liquid is entrained and incorporated into the thermal (figure 2a,b), a phenomenon referred to as turbulent entrainment. We observe the interaction between a large-scale engulfment, producing an entraining flow behind the thermal, and the development of small-scale turbulent fluctuations. As a result, turbulent entrainment progressively dilutes the fluid within the thermal, through stirring and mixing processes. In both experiments and simulations, we find that the radius of the thermal increases with depth in an approximately linear way, in agreement with the prediction of equation 18 (Appendix figure 12). All experiments (except the one at the lowest Reynolds) and simulations have close to an entrainment coefficient $\alpha$ of $0.18$, which is in agreement with previous experimental and numerical studies with an aspect ratio of the source equal to 1 (Lecoanet & Jeevanjee, 2019; Bond & Johari, 2010; Lai et al., 2015). While both experiments and simulations evolve similarly at large scale, in the close-up snapshots (figure 2c,d), we observe significant differences in the scale of the structures inside the turbulent thermal which are probably due to the difference of Schmidt number (see discussion in Sec. 5.4).
@@
Figure 2: (a,b) Evolution of a turbulent thermal for experiments with Reynolds number $Re$ equal to 3452 for an experiment and a simulation, respectively. On each image, the time is normalized by $t_{g}$. (c,d) Close-up snapshots of the last time-step for the experiment (a) and simulation (b).
4.2 Anatomy of a turbulent thermal
@@
A buoyant thermal is composed of a head, containing most of the initial buoyant fluid, followed by a turbulent wake. The head has an ellipsoidal shape flattened in the vertical direction. It consists of a vortex ring, the so-called cores of the thermal, and a peripheral region where the surrounding fluid is entrained in the thermal. The regions with high values of concentration are the undiluted cores (Blyth et al., 1988; Romps & Kuang, 2010).
@@
Although turbulent thermal structures are qualitatively well-defined, tracking their position, shape, and contour may be quantitatively challenging. In a miscible thermal, the entrainment of the surrounding fluid and the diffusion of buoyant fluid complicate the definition of its contour. Previous works implemented various methods, from an ellipsoidal shape defined by the horizontal and vertical average of the buoyancy field (Zhao et al., 2013; Lecoanet & Jeevanjee, 2019), to more sophisticated methods (Romps & Charn, 2015; McKim et al., 2020). Overall, the envelope of the thermal depends on a threshold value of the buoyancy field (temperature, concentration, or density).
@@
Figures 3 and 4 show the effect of the definition of the boundary of the thermal on its spatial extent and the concentration distribution. We investigate two different ways of defining the limits of the thermal: a tight threshold corresponding to a concentration equal to a fixed fraction $0.5$ of $\langle c\rangle$, the average of $c$ within the thermal defined by the mask, and a loose definition of the thermal limit corresponding to the same threshold with a dilation operation using the morphological structuring element. The threshold value is the same for all experiments and simulations; it is constant in terms of the value of $c/\langle c\rangle$, but time-dependent in terms of $c$. The reasons for this choice are explained in section 5.2. Both masks include concentration values of the head, cores, and wakes of the thermal.
@@
Figure 3a shows a turbulent thermal at $t/t_{g}=51.25$ and for $Re=4175$ and the two masks. The mask for a tight threshold is split into two parts (by a given $z/D$): the head of thermal and the wake. The experimental values of concentration are limited by the dynamic range of the camera (16-bit resolution), and they contain a maximum of 5 orders of magnitude in intensity variation. In figures 3b,c, we compare the histograms of the concentration values and the probability density functions of the different regions in the thermal. Most of the concentration values are below $10^{-3}$ and correspond to the background values outside the thermal (figure 3b). The histograms of the whole field view and both masks are very similar for concentration values above $10^{-3}$. The two parts (head and wake) of the tight mask show the same shape in the PDF (figure 3b), except there are no extreme values of concentration in the wake. These extreme values come for the most part from the cores inside the head, as already observed in previous experiments (Johari, 1992). Figure 3 shows that the loose mask includes values that are part of the background or surrounding fluid. Therefore, the shape of the PDF below $10^{-3}$ depends significantly on the image analysis, especially on the pixel size used in the morphological structuring element.
@@
Figure 3: (a) Snapshot of the concentration field at $t/t_{g}=51.25$ for one of the three experiments at $Re=4175$. Colored lines denote the mask applied on the snapshot to produce the PDF of the concentration in (b, c). (b) and (c) PDF of the concentration as a function of the concentration in log-log or semi-log, respectively. The dashed black line represents the PDF of the whole snapshot. The gray dotted line shows the PDF of the thermal using a loose threshold. The inside of the thermal comprises all pixels inside the cyan and red masks. Brown and cyan lines in (b,c) (and mask in (a)) denote the different parts of the thermal: wake and cores, respectively.
@@
In figure 4, we have applied the same threshold and mask to a simulation (with the same Reynolds number, at $t/t_{g}=31.6$). We reduced the size of the morphological structuring element to take into account the lower resolution of the simulation (figure 4a). While experiments have a camera-limited range of concentration values, the histogram of concentration values extends over 8 orders of magnitude in the simulation. For a concentration value above $10^{-2}$, the histograms of loose and tight masks are very similar. Most of the high values of concentration correspond to the cores of the thermal inside the head (figure 4b). In figure 4b, the PDF of the smallest values (between $10^{-8}$ and $10^{-5}$) decreases as $1/c$, which is a characteristic of the initial noise at $t/t_{g}=0$. The values of concentration between $10^{-5}$ and $10^{-2}$ are surrounding the thermal defined by the tight mask. The shape of the PDF for these values depends significantly on the image processing used to define the mask.
@@
Figure 4: (a) Snapshot of the concentration field at $t/t_{g}=31.6$ for one of the simulations at $Re=4175$. Colored lines denote the mask applied on the snapshot to produce the PDF of the concentration in (b, c). (b) and (c) PDF of the concentration as a function of the concentration in log-log or semi-log, respectively. The dashed black line represents the PDF of the whole snapshot. The gray dotted line shows the PDF of the thermal using a loose threshold. The inside of the thermal comprises all pixels inside the cyan and red masks. Brown and cyan lines in (b,c) (and mask in (a)) denote the different parts of the thermal: wake and cores, respectively.
@@
Based on figure 3 and 4, we have chosen to analyze our data using the tight threshold. Figure 3c and 4c show semi-log versions of the PDF of concentration, which highlights its shape for the values above $10^{-3}$. It suggests that the PDF of $c$ decreases exponentially at intermediate values of $c$, between the threshold and an upper value which is $\simeq 0.013$ and $\simeq 0.03$ in the experiment (figure 3b) and simulation (figure 4b) snapshots. The figures suggest that the deviations from the exponential PDF at high concentrations can be attributed to the cores. In contrast, the distribution in the wake seems to be well described by an exponential PDF.
5 Evolution of PDF and moments of concentration
@@
In this section, we investigate the moments and PDF of the concentration. We will examine the time dependence from the release of the buoyant fluid (for the simulations) to large distances and times after the release ($t/t_{g}>10$). We will also compare the differences and similarities between experiments and simulations that have been performed at the same Reynolds numbers but very different Péclet numbers. In what follows, the concentration is normalized by the initial concentration $c_{0}=c(t=0)$.
5.1 Moments of concentration
@@
Figure 5a,b shows the mean concentration $\langle c/c_{0}\rangle$, the maximum concentration against time, for all experiments and simulations. These are calculated from $t=0$ in the simulations, but we only focus on $t/t_{g}>1$. At larger $t/t_{g}$, the mean decreases following a power law close to $t^{-3/2}$, in agreement with the prediction of dimensional analysis (Eq. (22)). Since the concentration is averaged within each pixel, the absolute value of the maximum depends on the resolution of the concentration structures inside the thermal. In addition, the number of pixels covering the thermal is not constant, as the thermal radius increases as the square root of time. For example, in the simulations, there are five times more pixels per thermal radius at $t/t_{g}=30$ than at $t/t_{g}=0$. We quantified this effect by progressively reducing the resolution of a simulation over time, i.e., maintaining a constant number of pixels per thermal radius. From this, we conclude that the evolution of the maximum concentration observed in figure 5b is meaningful, i.e. the decrease of maximum concentration is not a geometrical artifact. For both experiments and simulations, the maximum concentration decreases more slowly than the mean concentration. The slopes of the maximum and mean concentrations seem similar between our experiments and simulations. However, the curves seem to be offset in time from one another. This might be due to the variation in the initial condition, which determines the value of $t_{g}$ and the time origin. We also found that these slopes agree reasonably well with the power-law fit from previous experiments for a similar Reynolds number (Zhao et al., 2013). Note that the fit from (Zhao et al., 2013) is obtained in a short time range. Figure 5c shows that the ratio between the maximum and the instantaneous concentration mean increases with time, which is consistent with the thermal cores evolving separately from the bulk of the thermals. The maximum concentration compared to the concentration’s mean in the simulations is smaller than in the experiments, which can be interpreted as the effect of a higher diffusion rate in the simulation.
@@
Figure 5: (a) Mean of concentration $c/c_{0}$ as a function of time $t/t_{g}$ for the average of three experiments at each $Re$ (solid lines) and the average of two simulations at each $Re$ (dashed lines). (b) Maximum concentration $c/c_{0}$ as a function of time $t/t_{g}$. In (a) and (b), the gray and black dashed lines show the power laws $t^{-3/2}$ and $t^{-0.84}$ as in Zhao et al. (2013). (c) The ratio between the maximum concentration and its mean as a function of time $t/t_{g}$. The red dashed line shows the ratio of the two previous power laws. Results of Zhao et al. (2013) (Z2013) span only between 40 and 100 $t/t_{g}$. Note that all power laws intend to show the slope, i.e., they have an arbitrary pre-factor.
@@
Figure 6 shows the $k$th moments of the concentration ($k=1$, 2, 4, 6, 8) over time. For each Reynolds number, we average three experiments and two simulations. The dashed grey lines show the predictions of the dimensional analysis (Eq. 22) for the time evolution of the moments, $\langle c^{k}\rangle\sim\left(\frac{c_{1}t}{t_{g}}\right)^{-\frac{3}{2}k}$. The values of the prefactors ($c_{1}=20$) and the implications for the shape of the PDF will be discussed in subsection 5.2. The time evolution of the moments is in good agreement with the predictions from dimensional analysis. Although the 8th moment of concentration decreases by 20 orders of magnitude over the investigated time interval, Eq. 22 still accurately predicts its evolution.
@@
Figure 6: Moment of concentration $c^{k},\,\text{with}\,k=1,2,4,6,8$ as a function of time $t/t_{g}$ for the average of three experiments at each $Re$ (solid lines) and the average of two simulations at each $Re$ (dashed lines). The gray dotted lines show the moments as predicted by dimensional analysis and the exponential shape of the PDF (Eq. 29).
5.2 PDFs of concentration

### v2

@@ III.2 Numerical simulations
We have conducted 3D simulations of thermals with the open-source pseudospectral code Dedalus [8, 7], using the setup described by [22]. The equations of conservation of momentum and concentration are solved under the Boussinesq approximation. Details about the simulations can be found in [22] as we follow the same approach. While experiments have a camera-limited range of concentration values (5 orders of magnitude), the precision of the simulations extends the value of the concentration field over 8 orders of magnitude. The thermal is initialized at $t=0$ by introducing a spherical concentration perturbation of diameter $d$ into an otherwise homogeneous fluid. The spatial domain extends horizontally over $10d$ in both the $x$ and $y$ directions, and vertically over $20D$ in the $z$ direction. It is respectively discretized in $256\times 256\times 512$ modes. We benchmark our simulations against a simulation with a higher resolution of $512\times 512\times 1024$, kindly provided by Daniel Lecoanet (see simulation at $Re=6300$ in [22]). At a given Reynolds number ($Re\approx 4631$), we find that the kinetic energy spectrum and PDF of concentration are similar between the two resolutions (see Fig. 12 in Supplementary Materials [15]).
@@
The numerical simulations use Reynolds numbers that match those of the experiments. In contrast, the Schmidt number $Sc$ is set to 1 in all numerical simulations (for computational resource reasons), resulting in a concentration field that diffuses much faster than in the experiments, where $Sc=2500$. Consequently, the Péclet number is also a factor $2500$ larger in the experiments than in the simulations. It will allow us to investigate the effect of the Schmidt number on concentration distributions.
@@
Note that turbulent thermals are at first order cylindrically symmetric. Then, we analyze two 2D planes orthogonal to each other from the 3D simulations, as it would be similar to the laser plane of the experiments. In Section V, we will show the evolution of concentration moments and PDF for the average of these two orthogonal planes and compare them to the experiments performed at the same Reynolds numbers.
IV Qualitative description of a turbulent thermal
IV.1 Macroscopic evolution
@@
Fig. 2 shows the evolution of the concentration field for a given $Re=3452$ for one experiment and a simulation. During its fall, the radius of the turbulent thermal increases as the surrounding liquid is entrained and incorporated into the thermal (Fig. 2a,b), a phenomenon referred to as turbulent entrainment. We observe the interaction between a large-scale engulfment, producing an entraining flow behind the thermal, and the development of small-scale turbulent fluctuations. As a result, turbulent entrainment progressively dilutes the fluid within the thermal, through stirring and mixing processes. In both experiments and simulations, we find that the radius of the thermal increases with depth, in agreement with the prediction of equation 18 (see Fig. 1 in Supplementary Materials [15]). Our simulations have an entrainment coefficient $\alpha$ close to $0.18$, as observed in previous numerical and experimental studies [22, 6, 17]. For the experiments, the entertainment coefficient is larger, especially for the smallest Reynolds number experiments ($\alpha\sim 0.4$ for the smallest $Re$, $\alpha\sim 0.25$ for the others $Re$), but this might be due to the difference in the release mechanism. The entrainment coefficient is very sensitive to the initial conditions, and is only accessible by averaging a large number of experiments [48].
Overall, both experiments and simulations evolve similarly at a large scale. The differences between the experiments and simulations on the depth of the thermals (at a given time) are due to the initial condition (release from a tube or from a sphere inside the surrounding fluid). While the Boussinesq approximation is used in the simulations, the experimental thermals experienced a large variation in density during the initial phase. In the close-up snapshots (Fig. 2c,d), we observe significant differences in the scale of the structures inside the turbulent thermal, which are probably due to the difference in Schmidt number (see discussion in Sec. V.4).
@@
Figure 2: (a,b) Evolution of a turbulent thermal for experiments with $Re=3452$ for an experiment ($Pe=8.64\times 10^{6}$) and a simulation $Pe=3452$, respectively. Note that the concentration scale changes between each time step for better visualization. On each image, the time is normalized by $t_{g}$. (c,d) Close-up snapshots of the last time step for the experiment (a) and simulation (b).
IV.2 Anatomy of a turbulent thermal
@@
A buoyant thermal is composed of a head, containing most of the initial buoyant fluid, followed by a turbulent wake. The head has an ellipsoidal shape flattened in the vertical direction. It consists of a vortex ring, the so-called cores of the thermal, and a peripheral region where the surrounding fluid is entrained in the thermal. The regions with high values of concentration are the undiluted cores [3, 33].
@@
Although turbulent thermal structures are qualitatively well-defined, tracking their position, shape, and contour may be quantitatively challenging. In a miscible thermal, the entrainment of the surrounding fluid and the diffusion of buoyant fluid complicate the definition of its contour. Previous works implemented various methods, from an ellipsoidal shape defined by the horizontal and vertical average of the buoyancy field [48] or using the velocity field [32, 22, 24]. Overall, the envelope of the thermal depends on a threshold value of the buoyancy field (temperature, concentration, or density).
@@
Figs. 3 and 4 show the effect of the definition of the boundary of the thermal on its spatial extent and the concentration distribution. We investigate two different ways of defining the limits of the thermal: a tight threshold corresponding to a concentration equal to a fixed fraction $0.5$ of $\langle c\rangle$, the average of $c$ within the thermal defined by the mask, and a loose definition of the thermal limit corresponding to the same threshold with a dilation operation using the nearest neighbors. The threshold value is the same for all experiments and simulations; it is constant in terms of the value of $c/\langle c\rangle$, but time-dependent in terms of $c$. The reasons for this choice are explained in section V.2. Both masks include concentration values of the head, cores, and wakes of the thermal. The wake is qualitatively determined by the material left behind the buoyant thermal by detrainement. In a turbulent thermal, detrainement is very sensitive to the Reynolds number [22], up to 10% of its initial volume for $Re>10^{4}$. We split the mask for a tight threshold, into two parts (by a given $z/D$): the head of thermal and the wake to identify their respective PDFs. We also qualitatively define the cores by using the extreme values with the criteria $c/c_{0}>(60\%\,\text{of}\,max(c/c_{0}))$ that here correspond to a changing slope in the PDF of $c/c_{0}$ (Fig. 3c).
@@
Fig. 3a shows a turbulent thermal with $Re=4175$ and the two masks. In Figs. 3b,c, we compare the probability density functions of the concentration values for the different regions in the thermal. Most of the concentration values are below $10^{-3}$ and correspond to the background values outside the thermal (Fig. 3b). The PDF of the whole field view and both masks are very similar for concentration values above $10^{-3}$. The two parts (head and wake) of the tight mask show the same shape in the PDF (Fig. 3b), except there are no extreme values of concentration in the wake. These extreme values come for the most part from the cores inside the head, as already observed in previous experiments [16]. Fig. 3 shows that the loose mask includes values that are part of the background or surrounding fluid. Therefore, the shape of the PDF below $10^{-3}$ depends significantly on the image analysis, especially the number of pixels considered within the thermal.
@@
Figure 3: (a) Snapshot of the concentration field at $t/t_{g}=51.25$ for one of the three experiments at $Re=4175$. Colored lines denote the mask applied to the snapshot to produce the PDF of the concentration in (b, c). (b) and (c) PDF of the concentration as a function of the concentration in log-log or semi-log, respectively. The dashed black line represents the PDF of the whole snapshot. The gray dotted line shows the PDF of the thermal using a loose threshold. The inside of the thermal comprises all pixels inside the cyan and red masks. Brown and cyan lines in (b,c) (and mask in (a)) denote the different parts of the thermal: wake and cores, respectively.
@@
In Fig. 4, we have applied the same threshold and mask to a simulation (with the same Reynolds number as in Fig. 3). For a concentration value above $10^{-2}$, the PDF values of loose and tight masks are very similar. Most of the high values of concentration correspond to the cores of the thermal inside the head (Fig. 4b). In Fig. 4b, the PDF of the smallest values (between $10^{-8}$ and $10^{-5}$) decreases as $1/c$, which is a characteristic of the initial perturbation at $t/t_{g}=0$. The values of concentration between $10^{-5}$ and $10^{-2}$ surround the thermal defined by the tight mask. The shape of the PDF for these values depends significantly on the image processing used to define the mask.
@@
Figure 4: (a) Snapshot of the concentration field at $t/t_{g}=31.6$ for one of the simulations at $Re=4175$. Colored lines denote the mask applied to the snapshot to produce the PDF of the concentration in (b, c). (b) and (c) PDF of the concentration as a function of the concentration in log-log or semi-log, respectively. The dashed black line represents the PDF of the whole snapshot. The gray dotted line shows the PDF of the thermal using a loose threshold. The inside of the thermal comprises all pixels inside the cyan and red masks. Brown and cyan lines in (b,c) (and mask in (a)) denote the different parts of the thermal: wake and cores, respectively.
@@
Based on Fig. 3 and 4, we have chosen to analyze our data using the tight threshold. Fig. 3c and 4c show semi-log versions of the PDF of concentration, which highlights its shape for the values above $10^{-3}$. It suggests that the PDF of $c$ decreases exponentially at intermediate values of $c$, between the threshold and an upper value which is $\simeq 0.013$ in the experiment (Fig. 3b) and $\simeq 0.03$ in the simulation (Fig. 4b) snapshots. The Figs. 3c and 4c suggest that the deviations from the exponential PDF at high concentrations can be attributed to the cores.
V Evolution of PDF and moments of concentration
@@
In this section, we investigate the moments and PDF of the concentration. We will examine the time dependence from the release of the buoyant fluid (for the simulations) to large distances and times after the release ($t/t_{g}>10$). We will also compare the differences and similarities between experiments and simulations that have been performed at the same Reynolds numbers but different Péclet numbers.
V.1 Moments of concentration
@@
Fig. 5a,b shows the mean concentration $\langle c/c_{0}\rangle$, the maximum concentration against time, for all experiments and simulations. These are calculated from $t=0$ in the simulations, but we only focus on $t/t_{g}>1$. At larger $t/t_{g}$, the mean decreases following a power law close to $t^{-3/2}$, in agreement with the prediction of dimensional analysis (Eq. (22)). We have checked the sensitivity of the evolution of the maximum concentration against the resolution. Since the concentration is averaged within each pixel, the absolute value of the maximum depends on the resolution of the concentration structures inside the thermal. In addition, the number of pixels covering the thermal is not constant, as the thermal radius increases as the square root of time (about 5 times more pixels per thermal radius at $t/t_{g}=30$ than at $t/t_{g}=0$). By quantifying the effect by progressively reducing the resolution of a simulation over time, i.e., maintaining a constant number of pixels per thermal radius, we conclude that the decrease of the maximum concentration is significant and not a geometrical artifact. For both experiments and simulations, the maximum concentration decreases more slowly than the mean concentration. The exponents of the maximum and mean concentrations are similar between our experiments and simulations. However, the experiments can be offset in time from one another. This is due to the inherent variability between each experiment in the initial condition, which determines the value of $t_{g}$ and the time origin. We also found that these slopes agree reasonably well with the power-law fit from previous experiments for a similar Reynolds number [48]. Note that the fit from [48] is obtained in a short time range, which can induce non-negligible error in the exponent. Fig. 5c shows that the ratio between the maximum and the instantaneous concentration mean increases with time, which is consistent with the thermal cores evolving separately from the bulk of the thermals. The maximum concentration compared to the concentration’s mean in the simulations is smaller than in the experiments, which can be interpreted as the effect of a higher diffusion rate in the simulation.
@@
Figure 5: (a) Mean of concentration $c/c_{0}$ as a function of time $t/t_{g}$ for the average of three experiments at each ReRe (solid lines) and the average of two simulations at each ReRe (dashed lines). (b) Maximum concentration $c/c_{0}$ as a function of time $t/t_{g}$. In (a) and (b), the gray and black dotted lines show the power laws $t^{-3/2}$ and $t^{-0.84}$ as in [48]. (c) The ratio between the maximum concentration and its mean as a function of time $t/t_{g}$. The red dashed line shows the ratio of the two previous power laws. Results of [48] (Z2013) span only between 40 and 100 $t/t_{g}$. Note that all power laws intend to show the slope, i.e., they have an arbitrary pre-factor.
@@
Fig. 6 shows the $k$th moments of the concentration ($k=1$, 2, 4, 6, 8) over time. For each Reynolds number, we average the three experiments and two orthogonal plans in each simulation. The dashed grey lines show the predictions of the dimensional analysis (Eq. 22) for the time evolution of the moments, $\langle c^{k}\rangle\sim\left(\frac{c_{1}t}{t_{g}}\right)^{-\frac{3}{2}k}$. The values of the prefactors ($c_{1}=20$) and the implications for the shape of the PDF will be discussed in subsection V.2. The time evolution of the moments is in good agreement with the predictions from dimensional analysis. Although the 8th moment of concentration decreases by 20 orders of magnitude over the investigated time interval, Eq. 22 still accurately predicts its evolution.
@@
Figure 6: Moment of concentration $c^{k},\,\text{with}\,k=1,2,4,6,8$ as a function of time $t/t_{g}$ for the average of three experiments at each ReRe (solid lines) and the average of two simulations at each ReRe (dashed lines). The gray dotted lines show the moments as predicted by dimensional analysis and the exponential shape of the PDF (Eq. 29).
V.2 PDFs of concentration